please help!
In a right triangle, find the length of the side not given
a=? b=1 c= sqr.root of 5
I'm totally stuck!
Thanks,

Answer by stanbon(75887)

In a right triangle, find the length of the side not given
a=? b=1 c= sqr.root of 5
-------------------------------
Since c^2 = a^2 + b^2
a^2 = c^2 - b^2
a^2 = (sqrt(5))^2 - 1^2
a^2 = 5 - 1
a = 2
==============
Cheers,
Stan H.